OpenStudy (jacob902):

Which of the following is the equation of a line that passes through the points (2,5) and (4,3)? A. y = -2x + 5 B. y = 2x + 5 C. y = -x + 7 D. y = x + 7

pooja195 (pooja195):

Find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (babynini):

Does that make sense? :) take your two points You have: (2, 5) which is (x1,y1) (4,3) which is (x2, y2)

pooja195 (pooja195):

The slope is not 2....try again.

OpenStudy (jacob902):

c?

OpenStudy (babynini):

remember that your equation looks like this: y=(slope)x+b So if you take your points and do what pooja said than you will find at least your slope! :)

OpenStudy (jacob902):

y = -x + 7?

pooja195 (pooja195):

That is correct

OpenStudy (babynini):

Mhm!! Exxcellent job.